OpenStudy (anonymous):

HELP! Please? So CONFUSED! Math hates me. :( Did I do it right? What is the sum of the first five terms of a geometric series with a1 = 6 and r = 1/3? Express your answer as an improper fraction in lowest terms. 6(1-1/3^5) 6(1-1/243) 6(242/243) s5 = --------- = ------- ------- 1 - 1/3 2/3 2/3 484/81 242 ------ = ---- 2/3 27 s5 = 242/27

OpenStudy (anonymous):

at the end I do 484/81 divided by 2/3 and get 242/27 but when I do 484/81 divided by 1-1/3 I get 457/81 Which one is correct?

OpenStudy (homeworksucks):

in your calculator is it

OpenStudy (homeworksucks):

1-(1/3)

OpenStudy (homeworksucks):

or 1-1/3

OpenStudy (anonymous):

it's supposed to be 1- 1/3, like that's what it is in the formula

OpenStudy (homeworksucks):

Yeah enter it in as 1-(1/3) not as 1-1/3 because your calculator thinks its (1-1)/3 I'm guessing

OpenStudy (homeworksucks):

Because take away 3/3 -1/3 and you get 2/3 which is right

OpenStudy (anonymous):

so the answer is 242/27?

OpenStudy (homeworksucks):

yup

OpenStudy (ranga):

Yes. 242/27. You can check your answer by this route also: 6 + 2 + 2/3 + 2/9 + 2/27 = 8 & 26/27 = 242/27